# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ II BD-RIS-Aided SISO: A Toy Example
$\bar{\rho}^{\mathsf{D-RIS}}=\mathbb{E}\left\{\left(\sum_{m=1}^{M}\left|[\mathbf{h}_{RI}]_{m}[\mathbf{h}_{IT}]_{m}\right|\right)^{2}\right\},$ (10)
which can be rewritten as
$\bar{\rho}^{\mathsf{D-RIS}}=\mathbb{E}\left\{\sum_{m=1}^{M}\left|[\mathbf{h}_{RI}]_{m}\right|^{2}\left|[\mathbf{h}_{IT}]_{m}\right|^{2}\right.\\ +\left.\sum_{m_{1}\neq m_{2}}\left|[\mathbf{h}_{RI}]_{m_{1}}\right|\left|[\mathbf{h}_{IT}]_{m_{1}}\right|\left|[\mathbf{h}_{RI}]_{m_{2}}\right|\left|[\mathbf{h}_{IT}]_{m_{2}}\right|\right\},$ (11)
where we expanded the square term. By noticing that $[\mathbf{h}_{RI}]_{m_{1}}$ and $[\mathbf{h}_{IT}]_{m_{2}}$ are independent $\forall m_{1},m_{2}$, and that $[\mathbf{h}_{RI}]_{m_{1}}$ and $[\mathbf{h}_{RI}]_{m_{2}}$ are independent if $m_{1}\neq m_{2}$, we obtain
@@
In Fig. 4, we report the average channel gain offered by D-RIS and BD-RIS under i.i.d. Rayleigh distributed channels. We compare the simulated channel gains, which are obtained by applying the proposed solutions to problems (4) and (7) and averaging over multiple channel realizations, with the theoretical channel gains derived in (15) and (13). As shown in Fig. 4, the simulation results confirm the accuracy of the closed-form scaling laws in (15) and (13). In the asymptotic regime where the number of RIS elements $M\to\infty$, the performance advantage of BD-RIS over D-RIS in a SISO system is quantified by the gain ratio
@@
implying that BD-RIS achieves approximately a 62% higher channel gain than D-RIS. Although this fundamental result highlights the clear advantage of BD-RIS in SISO systems, the benefits of BD-RIS become even more pronounced in multi-user and multi-antenna scenarios, as will be discussed in subsequent sections.
@@ III Modeling, Mode Analysis, and Architecture Design of BD-RIS
In this section, we introduce the modeling, architectures, and modes of BD-RIS, including the technical details, graphical illustrations, summaries and discussions. The reader is invited to consult the Appendix for a concise reminder of the background knowledge on multi-port network analysis.
III-A Modeling and Classification
@@
An $M$-element passive RIS is generally modeled as an $M$-antenna array connected to an $M$-port reconfigurable impedance network as illustrated in Fig. 5, which can be characterized by its impedance matrix $\mathbf{Z}_{I}\in\mathbb{C}^{M\times M}$, admittance matrix $\mathbf{Y}_{I}\in\mathbb{C}^{M\times M}$, or scattering matrix $\mathbf{\Theta}\in\mathbb{C}^{M\times M}$ [26, 35]. We invite the readers to consult the appendix for a reminder of the definition of impedance, admittance, and scattering matrices and the relationship between the three. These matrices are linked to each other by
@@
where $\mathbf{Y}_{I}=\mathbf{Z}_{I}^{-1}$, $Z_{0}$ denotes the reference impedance and $Y_{0}=Z_{0}^{-1}$ denotes the reference admittance used to calculate the scattering matrix. According to microwave network theory [34], for a passive reconfigurable impedance network, we have
@@
$\Re\{\mathbf{Z}_{I}\}=\mathbf{0}_{M\times M},~{}\Re\{\mathbf{Y}_{I}\}=\mathbf{0}_{M\times M}.$ (19)
Plugging (19) into (17), $\mathbf{\Theta}$ for a lossless reconfigurable impedance network is constrained by
@@
III-A.1 D-RIS and BD-RIS
Based on whether the three matrices, $\mathbf{Z}_{I}$, $\mathbf{Y}_{I}$, and $\mathbf{\Theta}$ are diagonal or not, we categorize RIS as D-RIS with diagonal matrices and BD-RIS with matrices not limited to being diagonal [25].
Remark 1: Here we give toy examples for a 4-element D-RIS with a diagonal scattering matrix $\mathbf{\Theta}^{\mathsf{D}}$ and a 4-element BD-RIS with a full scattering matrix $\mathbf{\Theta}^{\mathsf{BD}}$. Specifically in the lossless form of D-RIS, (20) indicates that $|[\mathbf{\Theta}^{\mathsf{D}}]_{1,1}|=\ldots=|[\mathbf{\Theta}^{\mathsf{D}}]_{4,4}|=1$, which is consistent with the constraint commonly considered in existing D-RIS literature [8, 9]. This shows that a lossless D-RIS can only support phase control of signals. However, for BD-RIS, (20) indicates that $\|[\mathbf{\Theta}^{\mathsf{BD}}]_{:,1}\|_{2}=\ldots=\|[\mathbf{\Theta}^{\mathsf{BD}}]_{:,4}\|_{2}=1$. Therefore, we have $|[\mathbf{\Theta}^{\mathsf{BD}}]_{m,n}|\leq 1$, $\forall m,n=1,\ldots,4$, implying that both amplitudes and phases of entries in $\mathbf{\Theta}^{\mathsf{BD}}$ are tunable even in the lossless case.
III-A.2 Reciprocal and Non-Reciprocal BD-RIS
@@ III-B.1 Reflecting Mode
III-B.2 Hybrid Transmitting and Reflecting Mode
In this mode, every two elements are placed back to back and interconnected with each other via reconfigurable admittance components, as illustrated in Fig. 7(b) [36]. With this arrangement, signals impinging on one sector of the RIS can be partially reflected toward the same sector and partially scattered toward the other sector, yielding a full-space coverage. This is also known as simultaneous transmitting and reflecting surface (STARS), STAR-RIS, or intelligent omni-surface (IOS) [37]. An extreme case of such a hybrid mode refers to the transmitting mode, which means the signals impinging on one sector of the surface are purely scattered toward the other sector.
III-B.3 Multi-Sector Mode
In this mode, every $L$, $L\geq 2$ elements are placed at the edge of an $L$-sided polygon and interconnected with each other via reconfigurable admittance components, as illustrated in Fig. 7(c) [38]. With this arrangement, signals impinging on one sector of RIS can be partially reflected toward the same sector and partially scattered toward the other sectors. More importantly, this arrangement allows each antenna to have a uni-directional radiation pattern covering only $\frac{1}{L}$ space to avoid overlapping among sectors, which provides higher gains than the hybrid mode. Consequently, multi-sector mode includes hybrid mode as a special case with $L=2$, while it goes beyond hybrid mode to achieve highly directional full-space coverage.
@@ III-B.3 Multi-Sector Mode
III-C Architecture Design: Reciprocal Architectures
In the $M$-port reciprocal reconfigurable impedance network, each port is connected to ground via its own reconfigurable admittance [^1] We use admittance parameter instead of impedance parameter to describe each reconfigurable component since admittance parameter describes the short-circuited property, that is how easily a circuit allows a current to flow. This allows us to have an admittance component with simply a zero admittance to describe that two elements are not connected. By contrast, the corresponding impedance component between the two elements will have an infinite value, which is not as straightforward as using admittance parameter. This will also simplify the mathematical modeling of various BD-RIS architectures, as will be detailed in the sequel. component, $Y_{m}$, and to port $m^{\prime}$, $m^{\prime}\neq m$, $\forall m,m^{\prime}\in\mathcal{M}$ via a reconfigurable admittance component $Y_{m,m^{\prime}}$, which satisfies $Y_{m,m^{\prime}}=Y_{m^{\prime},m}$ for $m^{\prime}>m$. Hence, the admittance matrix [^2] We use admittance matrix $\mathbf{Y}_{I}$ instead of impedance matrix $\mathbf{Z}_{I}$ to describe BD-RIS architectures since, according to (23), there is a linear mapping between the locations of nonzero entries in $\mathbf{Y}_{I}$ and the circuit topology of a given architecture. For example, when $Y_{m,m^{\prime}}=0$, $m\neq m^{\prime}$, indicating that ports $m$ and $m^{\prime}$ are not connected, the $(m,m^{\prime})$-th entry of $\mathbf{Y}_{I}$ is zero. By contrast, the corresponding impedance parameter-based descriptions have $Z_{m,m^{\prime}}=Y_{m,m^{\prime}}^{-1}$, $m\neq m^{\prime}$, $Z_{m}=Y_{m}^{-1}$ and $\mathbf{Z}_{I}=\mathbf{Y}_{I}^{-1}$, implying a nonlinear mapping between the structure of $\mathbf{Z}_{I}$ and certain circuit topologies. This means when $Y_{m,m^{\prime}}=0$, equivalently $Z_{m,m^{\prime}}=\infty$, the $(m,m^{\prime})$-th entry of $\mathbf{Z}_{I}$ could be neither zero nor $\infty$. $\mathbf{Y}_{I}$ of the reconfigurable impedance network is symmetric and each entry, $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ linking ports $m$ and $m^{\prime}$, can be calculated by making all other ports short-circuited [34]. Therefore, we have
$[\mathbf{Y}_{I}]_{m,m^{\prime}}=\begin{cases}-Y_{m,m^{\prime}},&m\neq m^{\prime},\\ Y_{m}+\sum_{k\neq m}Y_{m,k},&m=m^{\prime}.\end{cases}$ (23)
@@ III-C Architecture Design: Reciprocal Architectures
where $\varTheta_{m}=\frac{Y_{0}-Y_{m}}{Y_{0}+Y_{m}}$ following (17). The single-connected RIS is also referred to as D-RIS and has been widely studied and applied in wireless communication systems [8, 9].
III-C.2 Fully-Connected
To break through the strict mathematical constraint of $\mathbf{\Theta}$ in D-RIS and fully make use of the off-diagonal entries in $\mathbf{\Theta}$, a fully-connected architecture which could effectively improve the wave manipulation flexibility of RIS has been proposed in [26]. In this architecture, each port $m$ in the reconfigurable impedance network is connected to ground via $Y_{m}$ and to another port $m^{\prime}$, $m^{\prime}\neq m$, $\forall m,m^{\prime}\in\mathcal{M}$ via $Y_{m,m^{\prime}}$, which satisfies $Y_{m,m^{\prime}}=Y_{m^{\prime},m}$ for $m^{\prime}>m$. This indicates that the fully-connected architecture is realized by in total $\frac{M(M+1)}{2}$ reconfigurable admittance components. An illustrative example for a 4-element BD-RIS with fully-connected architecture is given in Fig. 8(b). Hence, $\mathbf{Y}_{I}$ is a full and symmetric matrix. By (17) and (19), for a lossless network, its impedance matrix $\mathbf{Z}_{I}$ is also a full and symmetric matrix, and the scattering matrix $\mathbf{\Theta}$ is a symmetric and unitary matrix, i.e.,

### v2

@@ II BD-RIS-Aided SISO: A Toy Example
$\bar{\rho}^{\mathsf{D-RIS}}$ $=\mathbb{E}\Big\{\Big(\sum_{m=1}^{M}\left|[\mathbf{h}_{RI}]_{m}[\mathbf{h}_{IT}]_{m}\right|\Big)^{2}\Big\}$ (6)
$=\mathbb{E}\Big\{\sum_{m=1}^{M}\left|[\mathbf{h}_{RI}]_{m}\right|^{2}\left|[\mathbf{h}_{IT}]_{m}\right|^{2}+\sum_{m_{1}\neq m_{2}}\left|[\mathbf{h}_{RI}]_{m_{1}}\right|$
$~~~~\times\left|[\mathbf{h}_{IT}]_{m_{1}}\right|\left|[\mathbf{h}_{RI}]_{m_{2}}\right|\left|[\mathbf{h}_{IT}]_{m_{2}}\right|\Big\},$
by expanding the square term. By noticing that $\mathbf{h}_{RI}]_{m_{1}}$ and $\mathbf{h}_{IT}]_{m_{2}}$ are independent $\forall m_{1},m_{2}$, and that $\mathbf{h}_{RI}]_{m_{1}}$ and $\mathbf{h}_{RI}]_{m_{2}}$ are independent if $m_{1}\neq m_{2}$, we obtain
@@
In Fig. 4, we report the average channel gain offered by D-RIS and BD-RIS under i.i.d. Rayleigh distributed channels. We compare the simulated channel gains, which are obtained by applying the proposed solutions detailed in Section IV-A and averaging over multiple channel realizations, with the theoretical channel gains derived in (10) and (8). As shown in Fig. 4, the simulation results confirm the accuracy of the closed-form scaling laws in (10) and (8). In the asymptotic regime where the number of RIS elements $M\to\infty$, the performance advantage of BD-RIS over D-RIS in a SISO system is quantified by the gain ratio
@@
implying that BD-RIS achieves approximately a 62% higher channel gain than D-RIS. Although this fundamental result highlights the clear advantage of BD-RIS in SISO systems, the benefits of BD-RIS become even more pronounced in multi-user and multi-antenna scenarios. For instance, BD-RIS can provide up to 75% sum-rate improvement over D-RIS with $M=128$, as will be shown in Section V.
@@ III Modeling, Mode Analysis, and Architecture Design of BD-RIS
In this section, we introduce the modeling, architectures, and modes of BD-RIS, including the technical details, graphical illustrations, summaries and discussions.
III-A Preliminaries on Multi-Port Network Analysis
As a foundation on the BD-RIS modeling, we start by concisely reviewing the background knowledge on multi-port network analysis, which is a simplified way to understand transmission properties compared to field analysis using Maxwell’s equations, and a powerful and useful technique for modeling and studying wireless systems [pozar2021microwave]. Specifically, each antenna in a wireless system is regarded as a port, whose behavior is characterized by its terminal voltage, current, incident wave, and reflected wave. This technique has been used to model and analyze MIMO systems [wallace2004mutual, wallace2004termination, morris2005network], where some hardware impairments, such as mutual coupling and antenna mismatching, can be explicitly captured based on physics-consistent assumptions. Due to the fact that RIS is essentially a multi-port network regarding each element as one port, the multi-port network theory is naturally suitable for modeling RIS. Therefore, recently, this technique has also been used to model and study RIS-aided wireless communication systems [shen2021, universal2024, gradoni2021end, nossek2024physically], by regarding transmit antennas, receive antennas, and RIS elements as a whole multi-port network. In this subsection, we briefly explain the concept of this technique, together with its three important parameters, namely, impedance, admittance, and scattering parameters, to lay the foundation for the modeling of BD-RIS [pozar2021microwave].
Figure 5: An arbitrary $M$-port microwave network.
III-A.1 Impedance, Admittance, and Scattering Parameters
Consider an arbitrary $M$-port microwave network, as depicted in Fig. 5. At the $m$-th port, the terminal voltage, current, incident wave, and reflected wave are given by $v_{m}\in\mathbb{C}$, $i_{m}\in\mathbb{C}$, $a_{m}\in\mathbb{C}$, and $b_{m}\in\mathbb{C}$. Define the terminal voltage vector $\mathbf{v}=[v_{1},\ldots,v_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, the current vector $\mathbf{i}=[i_{1},\ldots,i_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, the incident wave vector $\mathbf{a}=[a_{1},\ldots,a_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, and the reflected wave vector $\mathbf{b}=[b_{1},\ldots,b_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$ for the $M$-port network. These four vectors are related to each other by $\mathbf{v}=\mathbf{a}+\mathbf{b}$, $\mathbf{i}=\frac{\mathbf{a}-\mathbf{b}}{Z_{0}}$. The impedance matrix $\mathbf{Z}\in\mathbb{C}^{M\times M}$ of the $M$-port network then relates the terminal voltage vector and current vector by $\mathbf{v}=\mathbf{Z}\mathbf{i}$. In the impedance matrix $\mathbf{Z}$, each entry $\mathbf{Z}]_{m,n}$ is the impedance between ports $m$ and $n$ when all other ports are open-circuited, i.e., $\mathbf{Z}]_{m,n}=\frac{v_{m}}{i_{n}}\Big|_{i_{k}=0,\forall k\neq n}$. Similarly, the admittance matrix $\mathbf{Y}\in\mathbb{C}^{M\times M}$ of the $M$-port network relates the current vector and voltage vector by $\mathbf{i}=\mathbf{Y}\mathbf{v}$. Therefore, we have $\mathbf{Y}=\mathbf{Z}^{-1}$. In $\mathbf{Y}$, each entry $\mathbf{Y}]_{m,n}$ is the admittance between ports $m$ and $n$ when all other ports are short-circuited, i.e., $\mathbf{Y}]_{m,n}=\frac{i_{m}}{v_{n}}\Big|_{v_{k}=0,\forall k\neq n}$. The scattering matrix $\mathbf{S}\in\mathbb{C}^{M\times M}$ relates the incident wave vector and reflected wave vector by $\mathbf{b}=\mathbf{S}\mathbf{a}$. In $\mathbf{S}$, each entry $\mathbf{S}]_{m,n}$ is the scattering coefficient from port $n$ to port $m$ when all other ports are terminated in matched loads, i.e., $\mathbf{S}]_{m,n}=\frac{b_{m}}{a_{n}}\Big|_{a_{k}=0,\forall k\neq n}$. Based on the above illustration, the three parameters are related to each other by
$=(Y_{0}\mathbf{I}_{M}+\mathbf{Y})^{-1}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}),$
III-A.2 Reciprocal and Lossless Networks
$\mathbf{Z}=\mathbf{Z}^{\mathsf{T}},\mathbf{Y}=\mathbf{Y}^{\mathsf{T}},\mathbf{S}=\mathbf{S}^{\mathsf{T}}.$ (13)
$\Re\{\mathbf{Z}\}=\mathbf{0}_{M\times M},\Re\{\mathbf{Y}\}=\mathbf{0}_{M\times M},\mathbf{S}^{\mathsf{H}}\mathbf{S}=\mathbf{I}_{M}.$ (15)
III-B Modeling and Classification
@@
An $M$-element passive RIS is generally modeled as an $M$-antenna array connected to an $M$-port reconfigurable impedance network as illustrated in Fig. 6, which can be characterized by its impedance matrix $\mathbf{Z}_{I}\in\mathbb{C}^{M\times M}$, admittance matrix $\mathbf{Y}_{I}\in\mathbb{C}^{M\times M}$, or scattering matrix $\mathbf{\Theta}\in\mathbb{C}^{M\times M}$ [shen2021, universal2024]. These matrices are linked to each other by
@@
where $\mathbf{Y}_{I}=\mathbf{Z}_{I}^{-1}$. According to microwave network theory [pozar2021microwave], for a passive reconfigurable impedance network, we have
@@
III-B.1 D-RIS and BD-RIS
Based on whether the three matrices, $\mathbf{Z}_{I}$, $\mathbf{Y}_{I}$, and $\mathbf{\Theta}$ are diagonal or not, we categorize RIS as D-RIS with diagonal matrices and BD-RIS with matrices not limited to being diagonal [li2023reconfigurable].
Remark 1: Here we give toy examples for a 4-element D-RIS with a diagonal scattering matrix $\mathbf{\Theta}^{\mathsf{D}}$ and a 4-element BD-RIS with a full scattering matrix $\mathbf{\Theta}^{\mathsf{BD}}$. Specifically in the lossless form of D-RIS, (18) indicates that $|[\mathbf{\Theta}^{\mathsf{D}}]_{1,1}|=\ldots=|[\mathbf{\Theta}^{\mathsf{D}}]_{4,4}|=1$, which is consistent with the constraint commonly considered in existing D-RIS literature [di2020smart, wu2025intelligent]. This shows that a lossless D-RIS can only support phase control of signals. However, for BD-RIS, (18) indicates that $\|[\mathbf{\Theta}^{\mathsf{BD}}]_{:,1}\|_{2}=\ldots=\|[\mathbf{\Theta}^{\mathsf{BD}}]_{:,4}\|_{2}=1$. Therefore, we have $|[\mathbf{\Theta}^{\mathsf{BD}}]_{m,n}|\leq 1$, $\forall m,n=1,\ldots,4$, implying that both amplitudes and phases of entries in $\mathbf{\Theta}^{\mathsf{BD}}$ are tunable even in the lossless case.
III-B.2 Reciprocal and Non-Reciprocal BD-RIS
@@ III-C.1 Reflecting Mode
III-C.2 Hybrid Transmitting and Reflecting Mode
In this mode, every two elements are placed back to back and interconnected with each other via reconfigurable admittance components, as illustrated in Fig. 8(b) [li2022beyond]. With this arrangement, signals impinging on one sector of the RIS can be partially reflected toward the same sector and partially scattered toward the other sector, yielding a full-space coverage. This is also known as simultaneous transmitting and reflecting surface (STARS), STAR-RIS, or intelligent omni-surface (IOS) [zhang2022intelligent]. An extreme case of such a hybrid mode refers to the transmitting mode, which means the signals impinging on one sector of the surface are purely scattered toward the other sector.
III-C.3 Multi-Sector Mode
In this mode, every $L$, $L\geq 2$ elements are placed at the edge of an $L$-sided polygon and interconnected with each other via reconfigurable admittance components, as illustrated in Fig. 8(c) [li2023beyond]. With this arrangement, signals impinging on one sector of RIS can be partially reflected toward the same sector and partially scattered toward the other sectors. More importantly, this arrangement allows each antenna to have a uni-directional radiation pattern covering only $\frac{1}{L}$ space to avoid overlapping among sectors, which provides higher gains than the hybrid mode. Consequently, multi-sector mode includes hybrid mode as a special case with $L=2$, while it goes beyond hybrid mode to achieve highly directional full-space coverage.
@@ III-C.3 Multi-Sector Mode
III-D Architecture Design: Reciprocal Architectures
In the $M$-port reciprocal reconfigurable impedance network, each port is connected to ground via its own reconfigurable admittance [^1] We use admittance parameter instead of impedance parameter to describe each reconfigurable component since admittance parameter describes the short-circuited property, that is how easily a circuit allows a current to flow. This allows us to have an admittance component with simply a zero admittance to describe that two elements are not connected. By contrast, the corresponding impedance component between the two elements will have an infinite value, which is not as straightforward as using admittance parameter. This will also simplify the mathematical modeling of various BD-RIS architectures, as will be detailed in the sequel. component, $Y_{m}$, and to port $^{\prime}, mm$$^{\prime}\neq m$, $\forall m,m^{\prime}\in\mathcal{M}$ via a reconfigurable admittance component $Y_{m,m^{\prime}}$, which satisfies $Y_{m,m^{\prime}}=Y_{m^{\prime},m}$ for $m^{\prime}>m$. Hence, the admittance matrix [^2] We use admittance matrix $\mathbf{Y}_{I}$ instead of impedance matrix $\mathbf{Z}_{I}$ to describe BD-RIS architectures since, according to (21), there is a linear mapping between the locations of nonzero entries in $\mathbf{Y}_{I}$ and the circuit topology of a given architecture. For example, when $Y_{m,m^{\prime}}=0$, $m\neq m^{\prime}$, indicating that ports $m$ and $m^{\prime}$ are not connected, the $(m,m^{\prime})$-th entry of $\mathbf{Y}_{I}$ is zero. By contrast, the corresponding impedance parameter-based descriptions have $Z_{m,m^{\prime}}=Y_{m,m^{\prime}}^{-1}$, $m\neq m^{\prime}$, $Z_{m}=Y_{m}^{-1}$ and $\mathbf{Z}_{I}=\mathbf{Y}_{I}^{-1}$, implying a nonlinear mapping between the structure of $\mathbf{Z}_{I}$ and certain circuit topologies. This means when $Y_{m,m^{\prime}}=0$, equivalently $Z_{m,m^{\prime}}=\infty$, the $(m,m^{\prime})$-th entry of $\mathbf{Z}_{I}$ could be neither zero nor $\infty$. $\mathbf{Y}_{I}$ of the reconfigurable impedance network is symmetric and each entry, $\mathbf{Y}_{I}]_{m,m^{\prime}}$ linking ports $m$ and $m^{\prime}$, can be calculated by making all other ports short-circuited [pozar2021microwave]. Therefore, we have
$\mathbf{Y}_{I}]_{m,m^{\prime}}=\begin{cases}-Y_{m,m^{\prime}},&m\neq m^{\prime},\\ Y_{m}+\sum_{k\neq m}Y_{m,k},&m=m^{\prime}.\end{cases}$ (21)
@@ III-D Architecture Design: Reciprocal Architectures
where $\varTheta_{m}=\frac{Y_{0}-Y_{m}}{Y_{0}+Y_{m}}$ following (16). The single-connected RIS is also referred to as D-RIS with a diagonal matrix (3) in Section II and has been widely studied and applied in wireless communication systems [di2020smart, wu2025intelligent]. The corresponding scattering matrix $\mathbf{\Theta}$ is very often named as the phase-shift matrix since only the phase shift of each diagonal entry can be tuned.
III-D.2 Fully-Connected
To break through the strict mathematical constraint of $\mathbf{\Theta}$ in D-RIS and fully make use of the off-diagonal entries in $\mathbf{\Theta}$, a fully-connected architecture which could effectively improve the wave manipulation flexibility of RIS has been proposed in [shen2021]. In this architecture, each port $m$ in the reconfigurable impedance network is connected to ground via $Y_{m}$ and to another port $^{\prime}, mm$$^{\prime}\neq m$, $\forall m,m^{\prime}\in\mathcal{M}$ via $Y_{m,m^{\prime}}$, which satisfies $Y_{m,m^{\prime}}=Y_{m^{\prime},m}$ for $m^{\prime}>m$. This indicates that the fully-connected architecture is realized by in total $\frac{M(M+1)}{2}$ reconfigurable admittance components. An illustrative example for a 4-element BD-RIS with fully-connected architecture is given in Fig. 9(b). Hence, $\mathbf{Y}_{I}$ is a full and symmetric matrix. By (16), for a lossless network, its impedance matrix $\mathbf{Z}_{I}$ is also a full and symmetric matrix, and the scattering matrix $\mathbf{\Theta}$ is a symmetric and unitary matrix, i.e.,